A boy circles continuously through a distance of 1.0km in 5min.calculate the average speed
1
Expert's answer
2020-11-16T07:47:31-0500

"v=\frac{s}{t}\\v=\frac{1000}{5\cdot 60}=3.3\frac{m}{s}"